OpenStudy (4meisu):

An investment of P dollars made at the simple interest rate r for t years will grow to a total T= P +Prt. For an investment of $10,000 to grow to at least $12,000 in two years, what is the minimum interest rate at which it can be invested?

OpenStudy (anonymous):

i am also confused in this plz solve it all

OpenStudy (anonymous):

T= P +Prt 12000 = 10000 + 10000r(10) 2000 = 100000 r 2000/100000 = r 2/100 = r Rate is 2%

OpenStudy (4meisu):

Mm, no, it's not 2%, (I got the same thing) but the back of the book says 10% and I'm really confused as to how that's the answer.

OpenStudy (stormfire1):

Interest = Principal * rate * time so: 2000 = 10000 * r * (2) r = 10%.